A jet makes a trip to the maintenance facility and back. The trip there took 9 hours and the trip back took 6 hours. It averaged 330 km/h on the return trip. Find the average speed of the trip there.
Answer by mananth(16949)

Plane speed =x mph
Wind speed =y mph
against wind x-y 18.00 hours
with windx+y 9.00 hours

Distance = same= 864 miles
t=d/r
864/(x- y)=18.00
18(x-y)=864.00
18x-18y =864 ....................1
864/(x+y)= 9.00
9.00(x+y)= 864
9.00 x + 9.00 y = 864 ...............2
Multiply (1) by 1.00
Multiply (2) by 2.00
we get
18 x + -18 y = 864
18 x + 18 y = 1728
add the two
36 x = 2592
/ 36
x = 72 mph

plug value of x in (1)
18 x -18 y = 864
1296 -18 y = 864
-18 y = 864 -1296
-18 y = -432
y = 24 mph